$SPAD/src/input f02fjf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
n : Integer := 16;
k : Integer := 6
tol : DoubleFloat := 0.0001
novecs : Integer := 0
nrx : Integer := n
lwork : Integer := 86
lrwork : Integer := 1;
liwork : Integer := 1;
noits : Integer := 1000
m : Integer := 4;
x :Matrix SF:=new(nrx,k,0.0);
ifail : Integer := -1
a :Matrix FRAC INT:= new(n,n,0);
a(1,1) := 1;
a(1,2) := -1/4;
a(1,5) := -1/4;
for i in 2..4 repeat
  a(i,i-1) := -1/4
  a(i,i) := 1
  a(i,i+1) := -1/4
  a(i,i+4) := -1/4
for i in 5..n-4 repeat
  a(i,i-4) := -1/4
  a(i,i-1) := -1/4
  a(i,i) :=   1
  a(i,i+1) := -1/4
  a(i,i+4) := -1/4
for i in n-3..n-1 repeat
  a(i,i-4) := -1/4
  a(i,i-1) := -1/4
  a(i,i) :=   1
  a(i,i+1) := -1/4
a(16,16) := 1;
a(16,15) := -1/4;
a(16,12) := -1/4;

b:Matrix FRAC INT:= new(n,n,0);
b(1,1) := 1
b(1,2) := -1/2
for i in 2..n-1 repeat
  b(i,i-1) := -1/2
  b(i,i) := 1
  b(i,i+1) := -1/2
b(16,15) := -1/2
b(16,16) := 1
c : Matrix MachineFloat := (inverse (a))*b;
bb  := b :: Matrix MachineFloat
result:=f02fjf(n,k,tol,novecs,nrx,lwork,lrwork,liwork,m,noits,_
               x,ifail,bb :: ASP27('DOT),c :: ASP28('IMAGE))

———-

References^{[1] nothing}

5